OpenStudy (anonymous):

Use the graph of y=f(x) to estimate the value of f'(1). The given point is 1,3

OpenStudy (anonymous):

f(x) = y f(1) = 3 (1,3) f(2) = 6 (2,6) and so on.

OpenStudy (anonymous):

The answer in the back of the book is 2.

OpenStudy (anonymous):

How is the answer two if your answer is a graph?

OpenStudy (anonymous):

The graph is given because we are supposed to use it to derive.

OpenStudy (anonymous):

Oh,

OpenStudy (anonymous):

Use the graph of y=f(x) to estimate the value of f'(1). That's exactly what the question says.

OpenStudy (anonymous):

the f(x) at 1 is 3; therefore the point is 1,3